OpenStudy (anonymous):

how do I solve for x in the following, 3(4x+3)=3-4(x-1)

OpenStudy (aripotta):

do the distributive property first

OpenStudy (aripotta):

what do you get after distributing the 3 and the -4?

OpenStudy (aripotta):

@Jesusfelon you there?

OpenStudy (anonymous):

yes12x+9=3-4x-1

OpenStudy (aripotta):

ok, do 3 - (-1)

OpenStudy (anonymous):

I dont know where you got that

OpenStudy (aripotta):

on the right of the equation. i'm adding like terms, sorry i didn't explain that lol

OpenStudy (anonymous):

so do I subtract or multiply

OpenStudy (aripotta):

actually it would be 3 - 1 and you'd keep the -4

OpenStudy (aripotta):

-4x*

OpenStudy (aripotta):

wait wait wait. you distributed wrong!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I see it should be a -3 above

OpenStudy (aripotta):

here's what you should have gotten after distributing: 12x + 9 = 3 - 4x + 4

OpenStudy (anonymous):

no a -4?

OpenStudy (aripotta):

nope, because -4 x -1 = 4. when you multiply or divide two negative numbers, your product/quotient will be positive

OpenStudy (anonymous):

so that is a neg 4multiplied by -1?

OpenStudy (aripotta):

yes. we multiplied -4 and x to get -4x and -4 and -1 to get 4

OpenStudy (aripotta):

now you can add 3 and 4 and keep the -4x

OpenStudy (anonymous):

I get that confused, thank you

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

what do you after that?

OpenStudy (anonymous):

-4x from both sides

OpenStudy (aripotta):

no :/ i guess i'm not explaining this right. you should be getting 12x + 9 = -4x + 7. the 7 came from adding the 3 and 4.

OpenStudy (anonymous):

Yea You are doing great its me that is confused but we are at the same point.

OpenStudy (aripotta):

alright :) what do you do after that?

OpenStudy (anonymous):

I nee to get an x by its self

OpenStudy (aripotta):

right, but how are you going to do that? what do you think?

OpenStudy (anonymous):

add 4x to both sides

OpenStudy (aripotta):

that's what i did. so tell me the equation after that

OpenStudy (anonymous):

16x+9=7?

OpenStudy (aripotta):

yep :) and what's the equation after the next step?

OpenStudy (anonymous):

16x=-2?

OpenStudy (aripotta):

mhmm and now you get the answer. you'll have to simplify it though

OpenStudy (anonymous):

I have to / both sides by 16?

OpenStudy (aripotta):

yea, to get -2/16. what's the simplified answer?

OpenStudy (anonymous):

I don't know for sure, i'm not real good at fractions

OpenStudy (aripotta):

if you divide both the numerator and denominator by 2, what do you get?

OpenStudy (anonymous):

1/8?

OpenStudy (aripotta):

yes and no. remember to put the negative sign before it. since -2/2 = -1

OpenStudy (anonymous):

I told you I wasn't good at fractions,Lol

OpenStudy (aripotta):

lol it's alright. there's your answer :) i hope you understand what you did to get it

OpenStudy (anonymous):

Thank you and I think I understand it some what but I forget the rules sometimes.............ok a lot

OpenStudy (aripotta):

haha. if you ever need any help after this, feel free to tag me in your question if no one's helping you

OpenStudy (anonymous):

I appriciate that. I'm just starting a class so we might become good friends. Lol